# Deep Joint CSI Estimation-Feedback-Precoding for MU-MIMO OFDM Systems

### v1

@@ IV-A Separate Architecture
For the downlink precoding design processing, we consider two approaches: traditional precoding design and AI-based precoding design. In the traditional scheme, we employ the singular value decomposition (SVD) precoding method, which leverages the outputs, from $\dot{\mathbf{H}}_{d}(1)$ to $\dot{\mathbf{H}}_{d}(K)$, of the CSI reconstruction module for $K$ users. The SVD algorithm is applied to different subcarriers to obtain the right singular vectors and eigenvalues for the $K$ users. These right singular vectors are then used as the precoding vectors for transmitting the downlink data. Downlink power allocation is performed using the water-filling (WF) algorithm, which utilizes the eigenvalues and the downlink noise power $\sigma^{2}_{d}$ to determine the downlink transmission power for the $K$ users. For the AI-based precoding design, we use the network architecture ideas as [18] to jointly design the CSI feedback and precoding design networks. Both the CSI feedback module and the precoding design module share the same network architecture as JEFPNet to ensure a fair comparison. This approach is referred to as “JFPNet,” and its loss function is defined as maximizing the downlink spectral efficiency, using the ideally estimated CSI as the source for joint CSI feedback and precoding design.
IV-B Performance Comparison Among Different Architecture
TABLE I: Impact of CE and CSI FB accuracy on the performance of conventional SVD precoding algorithms
$\mathrm{SNR_{u}}$ Ideal CE & FB Ideal FB Ideal CE SVD & WF
-10 dB 14.405 6.687 3.923 3.874
-5 dB 14.405 6.687 4.372 4.275
0 dB 14.405 6.687 4.548 4.443
5 dB 14.405 6.687 4.611 4.505
10 dB 14.405 6.687 4.633 4.521
@@ IV-B Performance Comparison Among Different Architecture
Figure 7: Performance comparison of JEFPNet with traditional separate architecture and partial separate architecture for CSI estimation, feedback, and precoding.
Figure 8: Impact of different CE errors and overheads on the performance of separate and joint architectures.
@@
In this section, we compare the downlink spectral efficiencies of three different architectures: JEFPNet in a joint architecture, SVD precoding with WF power allocation in a separate architecture, and JFPNet in a partially separate architecture. The experiments are carried out with the $\mathrm{SNR_{ce}}$ in the CE phase and the $\mathrm{SNR_{d}}$ in the downlink data transmission phase fixed. The number of CE symbols ($L$) is set to 6, the feedback overhead ($Z$) is 16, and the number of users ($K$) is 2. By varying the SNR of the uplink feedback channel, i.e. $\mathrm{SNR_{u}}$, we evaluate the downlink spectral efficiency for each architecture. The experimental results are shown in Fig. 7. The curve labeled “JFPNet” depicts the JFPNet network architecture under ideal CE. The curve “SEFPNet” indicates the performance when the CE module is separately trained with JFPNet and tested together after concatenation. The “SVD_WF” curve represents the performance of the separate architecture, where the CE and CSI feedback modules are separately trained, and then concatenated for testing to generate reconstructed downlink CSI for precoding design with SVD and WF algorithms.
The experimental results show that, compared to traditional algorithms, represented by the “SVD_WF” curve in Fig. 7, the joint feedback and precoding design architecture consistently delivers performance gains. This is consistent with the findings of our previous work in [18]. Meanwhile, the proposed JEFPNet architecture delivering superior performance. At $\mathrm{SNR_{u}}=-10$ dB, JEFPNet achieves a performance gain of 4.97 bps/Hz over the traditional method depicted as the curve “SVD_WF,” and at $\mathrm{SNR_{u}}=10$ dB, the gain is 4.69 bps/Hz. Compared to the “SEFPNet”, which also account for CE and CSI feedback errors, JEFPNet achieves a performance gain of approximately 0.57 bps / Hz at $\mathrm{SNR_{u}}=-10$ dB and 0.60 bps/Hz at $\mathrm{SNR_{u}}=10$ dB.
Additionally, even when compared to “JFPNet,” which utilizes the ideal downlink full CSI as its source, the JEFPNet architecture still demonstrates a performance gain of 0.30 bps/Hz at $\mathrm{SNR_{u}}=-10$ dB. This improvement can be attributed to the joint estimation and feedback architecture employed by JEFPNet. Although both “SEFPNet” and “JFPNet” utilize CSI compression networks similar to JEFPNet, they compress the source using full CSI, while JEFPNet compresses the source using a lower-dimensional pilot signal instead. As a result, JEFPNet equivalently reduces the compression rate for the same output dimension. Meanwhile, with the assumption of limited transmission resources, directly extracting effective features from the received pilot signals with an end-to-end training manner with specific task objectives can enhance the efficiency of transmitting relevant semantic information, which in turn improves the downlink spectral efficiency.
@@
To investigate the performance loss of traditional algorithms, we compare the effects of CE errors and CSI feedback errors on traditional precoding design, with the results summarized in Table I. The experimental configurations in Table I align with those in Fig. 7. In Table I, “Ideal CE & FB” refers to the case that SVD precoding and the WF algorithm are used with ideal CE and CSI feedback. “Ideal FB” represents the case where erroneous CE is used with ideal CSI feedback, alongside the SVD precoding method. “Ideal CE” indicates the case where only the effect of CSI feedback on the traditional precoding algorithm is considered. The results in Table I show that the CSI feedback errors have a significant impact on the performance of the traditional algorithm. Specifically, when the feedback overhead $Z$ is 16 and the feedback channel quality $\mathrm{SNR_{u}}$ is -10 dB, CSI estimation errors lead to a performance degradation of 7.718 bps/Hz, while CSI feedback errors lead to a performance degradation of approximately 10.48 bps/Hz. Moreover, with the existence of CSI feedback errors, the addition of CE errors results in a performance loss of 0.049 bps/Hz. The smaller performance loss caused by CE error can be attributed to the large CE overhead ($L$) and the high $\mathrm{SNR_{ce}}$ of the downlink CE, both of which mitigate its impact, while the CSI feedback overhead remains relatively small. Nevertheless, the joint design of CSI feedback and precoding significantly reduces this performance loss, as demonstrated by the experimental results in Fig. 7 with curves “SEFPNet” and “JEFPNet”.
IV-C Advantage of End-to-End Multi-Module Joint Training
@@
JEFPNet is trained in an end-to-end manner that incorporates CE errors into the optimization of network parameters. In this subsection, we investigate the impact of estimation overhead and the channel quality during the CE phase on the final performance metrics (downlink spectral efficiency) of the network. The experimental results are presented in Fig. 8, where the three figures correspond to different channel qualities, i.e., $\mathrm{SNR_{ce}}\in\{-10\mathrm{dB},0\mathrm{dB},10\mathrm{dB}\}$. The impact of varying CE overheads on performance is also compared within each figure by considering two overhead values, $L=4$ and $L=6$, in our experiments.
@@
From the experimental results in Fig. 8, we observe that as feedback overhead and channel SNR decrease, the performance of both the separate and joint architectures deteriorates. However, the performance of JEFPNet under the joint architecture consistently outperforms than SEFPNet under the separate architecture. When $\mathrm{SNR_{ce}}$ is reduced from 10 dB to 0 dB, the performance of JEFPNet decreases by only 0.362 bps/Hz at $\mathrm{SNR_{u}}=-10$ dB for the $L=6$ case, while SEFPNet experiences a more drop of 0.431 bps/Hz under the same conditions. As $\mathrm{SNR_{ce}}$ further decreases from 10 dB to -10 dB, the difference between the two architectures becomes even more pronounced. The performance of JEFPNet decreases by 1.292 bps/Hz, while SEFPNet’s performance significantly decreases by 2.357 bps/Hz.
The reduction in CE overhead also impacts the downlink spectral efficiency. For $\mathrm{SNR_{ce}}=0$ dB, when the CE overhead $L$ is reduced from 6 to 4, the performance of JEFPNet degrades by approximately 0.1 bps/Hz in the range of $\mathrm{SNR_{u}}\in\left[{-10\mathrm{dB},10\mathrm{dB}}\right]$, whereas the performance of SEFPNet drops by about 0.36 bps/Hz.
The above phenomena can be attributed to the fact that SEFPNet uses ground-truth CSI for training. When the CE overhead is smaller, or the channel quality during estimation is poor, the CSI fed back during testing deviates more from the ground-truth CSI distribution, leading to degraded performance in the subsequent joint feedback and precoding design module.
IV-D Analysis of Storage and Computational Complexity Overhead
@@
In this subsection, we analyze the training parameters and floating point operations (FLOPs) of different architectures under specific configurations: the number of feedbacks $Z=16$, the number of subcarriers $N_{c}=96$, the CE overhead $L=6$, the pilot interval $g=4$, and the number of users $K=2$.
@@
Figure 9: Statistics on storage overhead and computational complexity for both sides of the communication system, considering different network architectures: LABEL:sub@UE_FLOPs the UE side; LABEL:sub@BS_FLOPs the BS side.
@@
Fig. LABEL:sub@UE_FLOPs illustrates the training parameters and FLOPs for different modules at each UE side. Bar plots represent the parameters and FLOPs of every module, while line plots show the total parameters and FLOPs for every vendor. In the JEFPNet architecture, the UE side consists only of a joint CE and CSI compression module, denoted as “Joint CE and CSI Compression”. In contrast, the partially separated SEFPNet and fully separated traditional architectures with SVD and WF algorithms, denoted as SVD_WF, implement CE and CSI compression as distinct modules, denoted as “CE” and “CSI Compression”, respectively. Statistical results reveal that the CE module in the separate architectures requires significantly more parameters and FLOPs. Additionally, the CSI compression module for feature extraction and CSI compression further increases the computational demand and parameter storage overhead. By comparison, the proposed JEFPNet architecture requires only 1.44% of the parameters and 0.0038% of the FLOPs compared to the separate architectures, making it more suitable for hardware-constrained UEs.
@@
Fig LABEL:sub@BS_FLOPs displays the parameters and FLOPs at BS. Since JEFPNet and SEFPNet share the same network architecture at the BS, their parameters and FLOPs, labeled as “Joint CSI Reconstruction and Precoding Design”, are identical. For the joint CSI reconstruction and precoding design module, since its network architecture depends on the number of users, both the parameters and FLOPs shown in Fig. LABEL:sub@BS_FLOPs correspond to the case of two users. For the traditional architecture, CSI reconstruction at the BS is implemented using $K$ decoders with shared parameters, labeled as “CSI Reconstruction”. Therefore, only one decoder’s parameters need to be stored and counted for statistics, which is shown on the left of Fig. LABEL:sub@BS_FLOPs. In the separate architecture, we employ SVD for precoding design and the WF algorithm for power allocation, collectively referred to as “SVD & WF”. Since the SVD and WF algorithms do not include trainable parameters, their parameter count is 0. For the FLOPs statistics shown on the right side of Fig. LABEL:sub@BS_FLOPs, the bar plot represents the FLOPs of every modular in the SVD_WF architecture, while the line plot corresponds to the total FLOPs at the BS vendor. As shown in Fig. LABEL:sub@BS_FLOPs, while the joint CSI reconstruction and precoding design in JEFPNet requires 225.6% more training parameters than the separate architecture, it reduces FLOPs by 52.12%. The increase in storage overhead is not an issue for the BS equipped with sufficient hardware capabilities. Comparatively, the reduction in FLOPs decreases CSI feedback and precoding design latency, thereby enhancing precoding reliability and mitigating the effects of CSI aging.
IV-E Scalability Tests
In this subsection, we experimentally verify the scalability of the proposed network under varying feedback overheads and numbers of users. First, we fix the number of users ($K$) to 2, the $\mathrm{SNR_{ce}}$ of the CE phase to 10 dB, and the CE overhead ($L$) to 6. We compare the downlink spectral efficiencies of different network architectures with single-user CSI feedback overhead ($Z$) values of 8, 16, and 24. The experimental results, shown in Fig. 10, indicate that as feedback overhead increases, the downlink spectral efficiency for both joint and separate architectures gradually improves. Notably, JEFPNet consistently outperforms JFPNet and SEFPNet. Specifically, the downlink spectral efficiency of JEFPNet exceeds that of SEFPNet by 0.382 bps/Hz when $Z=8$, 0.438 bps/Hz when $Z=16$, and 0.494 bps/Hz when $Z=24$, under $\mathrm{SNR_{ce}}=10$ dB.
Next, we fix $\mathrm{SNR_{ce}}$ to -10 dB, 0 dB, and 10 dB, the CE overhead to 6, and the feedback overhead to 32, while varying the number of users ($K$) between 2 and 3. The total transmit power $P$ in the data transmission phase is limited to 1, irrespective of the number of users. As shown in Fig. 11, increasing the number of users results in higher downlink spectral efficiency. Our proposed JEFPNet, implemented under the joint architecture, consistently outperforms SEFPNet, which operates under the separate architecture, particularly at low $\mathrm{SNR_{ce}}$. When $\mathrm{SNR_{u}}$ is fixed at -10 dB and $K=3$, the performance gain of JEFPNet over SEFPNet is 0.815 bps/Hz at $\mathrm{SNR_{ce}}=10$ dB, 0.720 bps/Hz at $\mathrm{SNR_{ce}}=0$ dB, and 1.76 bps/Hz at $\mathrm{SNR_{ce}}=-10$ dB. It is worth noting that, when $\mathrm{SNR_{u}}=0$ dB, the downlink spectral efficiency increases by 0.809 bps/Hz for three users compared to two users at $\mathrm{SNR_{ce}}=-10$ dB, by 0.955 bps/Hz at $\mathrm{SNR_{ce}}=0$ dB, and by 1.02 bps/Hz at $\mathrm{SNR_{ce}}=10$ dB. This improvement is attributed to the higher $\mathrm{SNR_{ce}}$ during the CE phase, which enables the acquisition of more accurate downlink CSI semantic information with the limitation of feedback overhead and CE overhead, facilitating the design of precoding vectors that effectively eliminate inter-user interference.

### v2

@@ IV-A Separate Architecture
For the downlink precoding design processing, we consider two approaches: traditional precoding design and AI-based precoding design. In the traditional scheme, we employ the singular value decomposition (SVD) and minimum mean squared error (MMSE) precoding methods, both of which utilize the outputs $\dot{\mathbf{H}}_{d}(1)$ to $\dot{\mathbf{H}}_{d}(K)$ from the CSI reconstruction module to design the precoding vectors for $K$ users. The SVD algorithm is applied to different subcarriers to obtain the right singular vectors and eigenvalues for the $K$ users. These right singular vectors are then used as the precoding vectors for transmitting the downlink data. Ideally, the precoding matrix obtained by the MMSE algorithm for the $n_{c}$-th subcarrier can be expressed as [MMSE_1, MMSE_2]
$\mathbf{V}_{n_{c}}=\mathbf{H}_{n_{c}}\left({\mathbf{H}_{n_{c}}^{H}\mathbf{H}_{n_{c}}+\frac{K\sigma_{d}^{2}}{P}\mathbf{I}}\right)^{-1},$ (25)
where $\mathbf{H}_{n_{c}}=\left[{\mathbf{h}_{d,n_{c}}^{T}(1),\cdots,\mathbf{h}_{d,n_{c}}^{T}(K)}\right]^{T}\in\mathbb{C}^{K\times N_{t}}$ is MU downlink channel matrix. Under non-ideal case, the design of $\mathbf{V}_{n_{c}}$ must rely on the reconstructed CSI, i.e., from $\dot{\mathbf{H}}_{d}(1)$ to $\dot{\mathbf{H}}_{d}(K)$. Downlink power allocation is performed using the water-filling (WF) algorithm, which utilizes the eigenvalues and the downlink noise power $\sigma^{2}_{d}$ to determine the downlink transmission power for the $K$ users. For the AI-based precoding design, we use the network architecture ideas as [JFPNet] to jointly design the CSI feedback and precoding design networks. Both the CSI feedback module and the precoding design module share the same network architecture as JEFPNet to ensure a fair comparison. This approach is referred to as “JFPNet,” and its loss function is defined as maximizing the downlink spectral efficiency, using the ideally estimated CSI as the source for joint CSI feedback and precoding design.
IV-B Performance Comparison Among Different Architectures
TABLE II: Impact of CE and CSI feedback errors on Traditional precoding performance
Algorithm $\bm{\mathrm{SNR_{u}}}$ Ideal Ideal FB Ideal CE Non-Ideal
SVD -10 dB 14.405 10.803 8.282 7.931
0 dB 14.405 10.803 9.581 9.119
10 dB 14.405 10.803 9.747 9.265
MMSE -10 dB 14.32 11.162 8.717 8.396
0 dB 14.405 11.162 9.925 9.514
10 dB 14.405 11.162 10.077 9.653
@@ IV-B Performance Comparison Among Different Architectures
Figure 7: Performance comparison of JEFPNet with traditional separate architecture and partial separate architecture for CE, feedback, and precoding design.
@@
In this section, we compare the downlink spectral efficiencies of three different architectures: JEFPNet in a joint architecture, SVD precoding and MMSE precoding with WF power allocation in a separate architecture, and JFPNet in a partially separate architecture. The experiments are carried out with the $\mathrm{SNR_{ce}}$ in the CE phase and the $\mathrm{SNR_{d}}$ in the downlink data transmission phase fixed. The number of CE symbols ($L$) is set to 6, the feedback overhead ($Z$) is 32, and the maximum number of users ($K_{\mathrm{max}}$) is 6. By varying the SNR of the uplink feedback channel, i.e. $\mathrm{SNR_{u}}$, we evaluate the downlink spectral efficiency for each architecture. We evaluate the network performance in Fig. 7 for the cases $K=2$ and $K=6$. The curve labeled “JFPNet” depicts the JFPNet network architecture under ideal CE. The curve “SEFPNet” indicates the performance when the CE module is separately trained with JFPNet and tested together after concatenation. The “SVD” and “MMSE” curves represent the performance of the separate architecture, where the CE and CSI feedback modules are separately trained, and then concatenated for testing to generate reconstructed downlink CSI for precoding design with SVD and MMSE algorithms.
The experimental results show that, for traditional algorithms, the errors introduced by CE and CSI feedback have a smaller impact on MMSE than on SVD. Compared with these traditional approaches, the proposed JEFPNet consistently achieves near-optimal performance, particularly at low $\mathrm{SNR_{u}}$. For example, with $K=2$ and $\mathrm{SNR_{u}}=-10$ dB, JEFPNet delivers performance gains of 1.42 bps/Hz over SVD and 0.96 bps/Hz over MMSE. Although MMSE slightly outperforms JEFPNet by 0.1 bps/Hz when $K=6$ and $\mathrm{SNR_{u}}=10$ dB, this does not undermine the performance advantage and effectiveness of the proposed network. This result can be attributed to the fact that JEFPNet must balance the overall average performance across different $\mathrm{SNR_{u}}$ and $K$ values simultaneously. Consequently, it prioritizes improving performance at low $\mathrm{SNR_{u}}$ and large $K$, as these scenarios exert a greater impact on the average performance. For instance, at $K=2$, JEFPNet achieves a gain of 0.956 bps/Hz over MMSE at $\mathrm{SNR_{u}}=-10$ dB, while the gain is only 0.029 bps/Hz at $\mathrm{SNR_{u}}=10$ dB. Furthermore, at $\mathrm{SNR_{u}}=-10$ dB and $K=6$, JEFPNet provides a gain of 2.67 bps/Hz compared to MMSE, and 0.959 bps/Hz at $K=2$. In the case of $K=2$, compared to the “SEFPNet”, which also accounts for CE and CSI feedback errors, JEFPNet achieves a performance gain of approximately 0.33 bps / Hz at $\mathrm{SNR_{u}}=-10$ dB and 0.17 bps/Hz at $\mathrm{SNR_{u}}=10$ dB.
To evaluate the significance of the pilot design module, we replace the AI-based pilot design module in JEFPNet with a traditional orthogonal discrete cosine transform (DFT)-based pilot design [CE_Methods], while keeping the rest of the architecture unchanged and continuing to train the network in an end-to-end manner. The results, shown by the “DFT_Pilot” curves in Fig. 7, reveal a substantial performance degradation compared with JEFPNet. Specifically, the performance decreases by 1.441̃.58 bps/Hz for $K=2$ and by approximately 3.053̃.39 bps/Hz for $K=6$. This degradation occurs because, under limited pilot resources, joint optimization of the pilot design module with other modules allows for the retention of more feature information relevant to the precoding task. However, the fixed pilot can’t effectively save feature information.
@@
To investigate the performance loss of traditional algorithms, we compare the effects of CE errors and CSI feedback errors on traditional precoding design, with the results summarized in Table II. The experimental configurations in Table II are identical to those in Fig. 7, except that we have set the evaluated $K$ to be 2. In Table II, “Ideal” refers to the case that traditional precoding and the WF algorithm are used with ideal CE and CSI feedback. “Ideal FB” represents the case where erroneous CE is used with ideal CSI feedback, alongside the traditional precoding method. “Ideal CE” indicates the case where only the effect of CSI feedback on the traditional precoding algorithm is considered. “Non-Ideal” denotes that errors from CE and CSI feedback are taken into account when employing traditional precoding algorithms. The results in Table II show that the CSI feedback errors have a significant impact on the performance of the traditional algorithm. Specifically, when the feedback overhead $Z$ is 32 and the feedback channel quality $\mathrm{SNR_{u}}$ is -10 dB, for the SVD precoding algorithm, CE errors lead to a performance degradation of 3.602 bps/Hz, while CSI feedback errors lead to a performance degradation of approximately 6.123 bps/Hz. Moreover, with the existence of CSI feedback errors, the addition of CE errors results in a performance loss of 0.351 bps/Hz. The smaller performance loss caused by CE error can be attributed to the large CE overhead ($L$) and the high $\mathrm{SNR_{ce}}$ of the downlink CE, both of which mitigate its impact, while the CSI feedback overhead remains relatively small. Nevertheless, the joint design of CSI feedback and precoding significantly reduces this performance loss, as demonstrated by the experimental results in Fig. 7 with curves “SEFPNet” and “JEFPNet”. For the MMSE precoding algorithm, under the same conditions, although the performance is 0.085 bps/Hz lower than that of the SVD algorithm in the “Ideal” case, the performance degradation is limited to 5.924 bps/Hz with the errors, which remains 0.46 bps/Hz higher than that of the SVD algorithm.
Figure 8: Network components. A direct connection between the regions corresponding to a specific module indicates that this module is assumed to be ideal. “Transformer” denotes the precoding design module in JEFPNet, which is trained separately; “Joint PD & CE” denotes the joint pilot design and CE network in SEFPNet; and “DJSCC FB” denotes the CSI feedback network based on the DJSCC architecture.
Figure 9: Ablation experiment.
To assess the impact of each module on system performance, we examine the effect of introducing different AI modules relative to the ideal case under the conditions $L=6$, $Z=32$, $\mathrm{SNR}_{u}=-10$ dB, $\mathrm{SNR}_{\mathrm{ce}}\in\{10~\text{dB},0~\text{dB}\}$. Six scenarios are considered, as illustrated in Fig. 8. The experimental results are shown in the Fig. 9
By comparing the performance of “CE/Ideal/Precoder” and “JEFPNet” at $\mathrm{SNR}_{\mathrm{CE}}=10$ dB and 0 dB, we observe that CE errors degrade the performance of the separated architecture more severely than that of the joint architecture. Specifically, at $K=6$ and $\mathrm{SNR}_{\mathrm{ce}}=0$ dB, CE error causes a performance loss of 11.76 bps/Hz, feedback error causes a loss of 12.42 bps/Hz, and their combined effect results in a loss of 14.70 bps/Hz. Compared with the separated architecture “CE/DJSCC/Precoder,” joint feedback and precoding design improves performance by 2.65 bps/Hz, whereas joint CE-feedback-precoding design achieves an improvement of 4.71 bps/Hz.
IV-C Adaptive Testing of the Number of Users
Figure 10: Adaptive Testing of the Number of Users. $K_{m}ax$ is setting 6 for architectures with user-number adaptability.
To demonstrate the effectiveness of the proposed architecture, we evaluate the performance of JEFPNet under both a fixed and a variable number of users, referred to as “JEFPNet (Transformer)” and “De-Adaptive (Transformer)”, respectively. At the same time, we replace the proposed precoding network in JEFPNet (Fig. LABEL:sub@PrecodingDesign_Transformer) with a user-adaptive precoding network with user-adaptive machine in [MU_Adaptive] (Fig. LABEL:sub@PrecodingDesign_DNN), referred to as “JEFPNet (DNN)”. Notably, “JEFPNet (Transformer)” requires 258,300 learnable parameters on the BS side, whereas “De-Adaptive (Transformer)” must store separate models for each user configuration, each comprising 258,300 parameters. The “JEFPNet (DNN)” architecture requires 273,148 learnable parameters. The remaining parameter settings are provided in the caption of Fig. 10. Experimental results demonstrate that “JEFPNet (Transformer)” outperforms “JEFPNet (DNN)”, particularly as the number of users increases. Specifically, at $\mathrm{SNR}_{u}=10$ dB and $K=6$, “JEFPNet (Transformer) achieves a performance gain of 2.78 bps/Hz. Furthermore, the performance of “JEFPNet (Transformer)” with user-number adaptability is comparable to that of “De-Adaptive (Transformer)”, with only a 0.11 bps/Hz performance gap under the same conditions.
By comparing “JEFPNet (Transformer)” and “De-Adaptive (Transformer),” we demonstrate that the proposed network achieves performance comparable to that of a network trained with a fixed number of users, while reducing network storage overhead by a factor of $1/K_{\max}$. Furthermore, the user-number adaptability design of JEFPNet eliminates the need for model switching when the number of users changes, making it particularly suitable for two-sided architectures where encoder models must be distributed to UEs from BS.
IV-D Evaluating the Validity of Joint CE within the DJSCC Framework
Figure 11: Impact of different CE errors on the performance of separate and joint architectures.
@@ IV-D Evaluating the Validity of Joint CE within the DJSCC Framework
JEFPNet is trained in an end-to-end manner that incorporates CE errors into the optimization of network parameters. In this subsection, we investigate the impact of estimation overhead and the channel quality during the CE phase on the final performance metrics (downlink spectral efficiency) of the network. In the experiment corresponding to Fig. 11, we fix $L$, set $K_{\max}=6$, $Z=32$, and $L=6$, and evaluate the impact of CE error rate on performance. Specifically, we compare JEFPNet with a joint architecture and SEFPNet with a separate architecture under $K\in\{2,6\}$ by varying $\mathrm{SNR_{ce}}$ at 10-10, 0, and $10$ dB. In the experiment corresponding to Fig. 12, we fix $\mathrm{SNR_{ce}}=10$ dB and $\mathrm{SNR_{u}}=-10$ dB, set $L\in\{2,4,6,16\}$, and keep other parameters consistent with those in Fig. 11. We then assess the impact of pilot length, which is also called CE overhead in this paper, on the performance of JEFPNet and SEFPNet under $K\in\{2,4,6\}$.
Figure 12: Impact of different CE overheads on the performance of separate and joint architectures.
@@
From the experimental results in Fig. 11, we observe that as the channel SNR decreases, the performance of both the separate and joint architectures deteriorates. However, the performance of JEFPNet under the joint architecture consistently outperforms SEFPNet under the separate architecture. When $\mathrm{SNR_{ce}}$ is reduced from 10 dB to 0 dB, the performance of JEFPNet decreases by only 0.193 bps/Hz at $\mathrm{SNR_{u}}=-10$ dB for the $L=6$ case, while SEFPNet experiences a greater drop of 0.894 bps/Hz under the same conditions. As $\mathrm{SNR_{ce}}$ further decreases from 10 dB to -10 dB, the difference between the two architectures becomes even more pronounced. The performance of JEFPNet decreases by 1.74 bps/Hz, while SEFPNet’s performance significantly decreases by 3.39 bps/Hz. The above phenomena can be attributed to the fact that SEFPNet uses ground-truth CSI for training. When the channel quality during estimation is poor, the CSI fed back during testing deviates more from the ground-truth CSI distribution, leading to degraded performance in the subsequent joint feedback and precoding design module.
We validate the effectiveness of CE within the DJSCC framework by evaluating SEFPNet with a separate CE design and JEFPNet with a joint CE and CSI feedback design. Increasing the CE overhead $L$ improves downlink spectral efficiency of JEFPNet and SEFPNet, as shown in Fig. 12. This trend arises because a larger $L$ enables JEFPNet to extract richer semantic information from the pilot signals received at the UE, while allowing SEFPNet to achieve higher CE accuracy and thereby decrease the impact of estimation errors. Specifically, JEFPNet outperforms JFPNet by 0.495 bps/Hz for $L=16$, $K=6$, and by 0.335 bps/Hz for $K=2$. Furthermore, as $L$ increases, the performance gap between JEFPNet and SEFPNet gradually narrows. For example, when $K=6$ under the same conditions, the gap is $0.419$ bps/Hz at $L=4$ and decreases to $0.320$ bps/Hz at $L=16$.
IV-E Analysis of Storage and Computational Complexity Overhead
@@
In this subsection, we analyze the training parameters and floating point operations (FLOPs) of different architectures under specific configurations: the number of feedbacks $Z=32$, the number of subcarriers $N_{c}=96$, the CE overhead $L=6$, the pilot interval $g=4$, and the number of users $K_{\max}=6$.
@@
Figure 13: Statistics on storage overhead and computational complexity for both sides of the communication system with $K=2$, considering different network architectures: LABEL:sub@UE_FLOPs the UE side; LABEL:sub@BS_FLOPs the BS side.
@@
Fig. LABEL:sub@UE_FLOPs illustrates the training parameters and FLOPs for different modules at each UE side. Bar plots represent the parameters and FLOPs of every module, while line plots show the total parameters and FLOPs for every vendor. In the JEFPNet architecture, the UE side consists only of a joint CE and CSI compression module, denoted as “Joint CE and CSI Compression”. In contrast, the partially separated SEFPNet and fully separated traditional architectures with MMSE and WF algorithms, denoted as MMSE, implement CE and CSI compression as distinct modules, denoted as “CE” and “CSI Compression”, respectively. Statistical results reveal that the CE module in the separate architectures requires significantly more parameters and FLOPs. Additionally, the CSI compression module for feature extraction and CSI compression further increases the computational demand and parameter storage overhead. By comparison, the proposed JEFPNet architecture requires only 2.63% of the parameters and 0.019% of the FLOPs compared to the separate architectures, making it more suitable for hardware-constrained UEs.
@@
Fig LABEL:sub@BS_FLOPs displays the parameters and FLOPs at BS with $K=2$. Since JEFPNet and SEFPNet share the same network architecture at the BS, their parameters and FLOPs, labeled as “Joint CSI Reconstruction and Precoding Design”, are identical. For the joint CSI reconstruction and precoding design module, since its network architecture depends on the number of users, both the parameters and FLOPs shown in Fig. LABEL:sub@BS_FLOPs correspond to the case of two users. For the traditional architecture, CSI reconstruction at the BS is implemented using $K$ decoders with shared parameters, labeled as “CSI Reconstruction”. Therefore, only one decoder’s parameters need to be stored and counted for statistics, which is shown on the left of Fig. LABEL:sub@BS_FLOPs. In the separate architecture, we employ MMSE for precoding design and the WF algorithm for power allocation, collectively referred to as “MMSE”. Since the MMSE and WF algorithms do not include trainable parameters, their parameter count is 0. For the FLOPs statistics shown on the right side of Fig. LABEL:sub@BS_FLOPs, the bar plot represents the FLOPs of every modular in the MMSE architecture, while the line plot corresponds to the total FLOPs at the BS vendor. As shown in Fig. LABEL:sub@BS_FLOPs, to avoid the complexity of model monitoring and switching at the BS, the proposed JEFPNet introduces a user-number-adaptive design. Consequently, in the two-user case, JEFPNet requires more training parameters and incurs higher FLOPs on the BS side due to the larger $K_{\max}$. However, without user-number adaptability design, when $K_{\max}=6$, the BS must store multiple models corresponding to different user numbers. Specifically, for $K=2$ to $6$, the non-adaptive network requires five separate models with a total parameter count of 1.2912M, whereas the adaptive design reduces this requirement to only 20% of that amount. Moreover, the non-adaptive network also necessitates training different encoder networks, leading to additional model delivery or storage overheads for the UE side.